• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ulmer

    Log entry time 16:05:45 on June15,2002

    Entry number 86102

    keyword=Shift summary

    Start shift with Q2D_c40 with roughly 65 uA beam
    We may have to live with 65 uA until Monday.

    09:08 Finish Q2D_c40
    09:18 Finish Q2D_c40 efficiency run
    09:20 Setting up for Q2D_cdN; run trigsetup for Q2D_d10
    09:45 MCC delivers 75 uA
    10:00 Start Q2D_cdN
    DAQ problems - reboot ROC14
    10:42 Finish Q2D_cdN
    10:55 MCC says they can now only deliver 50 uA of stable beam;
    they will try to get higher currents though.
    11:05 Start Q2D_d10
    12:12 End Q2D_d10
    12:20 As MCC finally gets us 85 uA, we need to take controlled access since
    Right arm angle camera does not respond.
    12:45 Finished access and setup for Q2D_d20 point.
    13:13 Begin Q2D_d20 point. Very poor beam.
    14:28 Beam went away ~ 20 min. ago - stop run and wait ...

    Still no stable beam at shift's end.